Serena and Peter leave their houses that are 5 miles apart and travel in opposite directions away from each other. Peter travels s mph and Serena travels 15 mph faster than Peter. i need all the following answered .
how many miles did peter travel in 2 hours?
how many miles did Serina travel in 2 hours?
how far apart are peter and Serina?
what is peters speed what is serinas speed?
Serina and Pete are 215 miles apart after 2 hours use this info and write an equation

Answer by math_tutor2020(3835)


s = Peter's speed
s+15 = Serena's speed since she is 15 mph faster

Calculate the distance Peter travels after 2 hours
distance = rate*time
distance = s*2
distance = 2s

Do the same for Serena
distance = rate*time
distance = (s+15)*2
distance = 2s+30

Add the distances and add the gap of 5 miles as well
(2s)+(2s+30)+(5) = 4s+35

After 2 hours, the two people are 4s+35 miles apart
Set this equal to the 215 mentioned and we get the following
4s+35 = 215
4s = 215-35
4s = 180
s = 180/4
s = 45

Peter is traveling at a rate of 45 mph
Serena is traveling at a rate of s+15 = 45+15 = 60 mph

-----------------------------------------------------------------

Question: How many miles did Peter travel in 2 hours?
Answer: 90 miles

Reason:
distance = rate*time = 45*2 = 90

-----------------------------------------------------------------

Question: How many miles did Serena travel in 2 hours?
Answer: 120 miles

Reason:
Distance = rate*time = 60*2 = 120

-----------------------------------------------------------------

Question: How far apart are Peter and Serena?
Answer: 215 miles

Reason:
Stated in the instructions

-----------------------------------------------------------------

Question: What is Peters speed what is Serena's speed?
Answer:
Peter's speed = 45 mph
Serena's speed = 60 mph

Reason:
Calculated earlier.